OpenStudy (anonymous):

Solve the equation 9^x+1 - 19 x 3^x + 2 = 0 giving 3 significant figures in your answer where appropriate.

terenzreignz (terenzreignz):

Is it like this \[\large 9^{x+1}-19x \cdot 3^{x} + 2=0\]

OpenStudy (anonymous):

19 multiply 3 power x

terenzreignz (terenzreignz):

Oh... \[\large \large 9^{x+1}-19 \cdot 3^{x} + 2=0\] this then?

OpenStudy (anonymous):

yea

terenzreignz (terenzreignz):

Okay, @Sgstudent , let's play a game. First, we let \[\large u = 3^x\] So we can write \[\large \large 9^{x+1}-19 \cdot 3^{x} + 2=9\cdot9^{x}-19\cdot3^{x}+2=9\cdot(3^x)^{2}-19(3^x)+2=9u^2 -19u+2\] \[\large 9u^2 -19u+2=0\] Can you do it now?

OpenStudy (anonymous):

oh lol i can do it now.

OpenStudy (anonymous):

How did you get (3^x)2 ??

terenzreignz (terenzreignz):

Laws of exponents. \[\huge (a^m)^n = a^{mn}\]

OpenStudy (anonymous):

From which a ^mn

terenzreignz (terenzreignz):

Okay, step by step, then :D \[\large 9^x = (3^2)^x = 3^{2x}=3^{x\cdot2}=(3^x)^2\]

OpenStudy (anonymous):

That's what i'm looking for thank you !

terenzreignz (terenzreignz):

No sweat :)

OpenStudy (anonymous):

where are you from?